OpenStudy (sonyalee77):

Two dice are rolled. Find the probability that the sum of the numbers on the two dice is 4.

OpenStudy (sonyalee77):

@Directrix

Directrix (directrix):

There are 6 possible outcomes for the first die and 6 outcomes for the second die. (Die is the singular of dice.) By the Fundamental Principle of Counting, that gives 6*6 = 36 possibilities of outcomes for the pair of dice.

Directrix (directrix):

Your task is to look at the attached chart above and count the number of throws that have spots that sum to 4. So , do that now and then we will get the probability.

Directrix (directrix):

@Sonyalee77 How many sums of 4 did you count? Let's knock out this problem.

OpenStudy (sonyalee77):

12

Directrix (directrix):

Would you write down the first six of them, please?

OpenStudy (sonyalee77):

IM SO CONFUSED

Directrix (directrix):

Look on the attached chart. There are only three ways to get a sum of four.

OpenStudy (sonyalee77):

OH I just counted all of the pairs with 4 in it

Directrix (directrix):

So when you got 12 sums of 4, I think you just looked for a four and counted it. That is not what we are doing. We are looking for sums of four. So, if you counted (4,4) then that is a 4 on one die and another 4 on the other. That adds to 8. We don't want that. We want a sum of four.

Directrix (directrix):

So, the probability of a 4 on a single toss of a pair of cubical dice is 3/36 which is 1/12. P(sum of 3) = 3/36 = 1/12.

OpenStudy (sonyalee77):

OOOOOOOHHHH Thank you so much